21 lbs of a metal is 29 percent nickel. it is a combination of metal A which is 35 percent nickel and metal b which is 25 percent nickel. how many pounds is metal A and how many pounds is metal B? The answer should be 8.4 and 12.6 but I don't know why
Answer by ankor@dixie-net.com(22740)

21 lbs of a metal is 29 percent nickel.
it is a combination of metal A which is 35 percent nickel and metal b
which is 25 percent nickel.
how many pounds is metal A and how many pounds is metal B?
:
Write an equation for each statement;
the total weight equation
a + b = 21
b = (21-a); use this form for substitution
:
the mixture equation in decimal form
.35a + .25b = .29(21)
Substitute (21-a) for b
.35a + .25(21-a) = .29(21)
do the math
.35a + 5.25 - .25a = 6.09
.35a - .25a = 6.09 - 5.25
.10a = .84
a = .84/.1
a = 8.4 lb of A mixture
then
b = 21 - 8.4
b = 12.6 lb of B mixture
:
:
;
See if that checks out
.35(8.4) + .25(12.6) =
2.94 + 3.15 = 6.09